# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ A.1 Properties of the Wasserstein Distance
Let $p,q\in{\mathcal{P}}({\mathbb{R}}^{n})$ be joint distributions that are independent over each dimension, i.e., $p=\prod_{i=1}^{n}{p^{{(i)}}}$ and $q=\prod_{i=1}^{n}{q^{{(i)}}}$, where ${p^{{(i)}}}$ and ${q^{{(i)}}}$ are the marginal distributions of $p$ and $q$ for the $i$-th dimension, respectively. Then, we have that:
${\mathbb{W}}_{2}^{2}(p,q)=\sum_{i=1}^{n}{\mathbb{W}}_{2}^{2}({p^{{(i)}}},{q^{{(i)}}}).$
Proof  Due to $p$ and $q$ being independent over each dimension, we have:
${\mathbb{W}}_{2}^{2}(p,q)$ $=\inf_{\gamma\in\Gamma(p,q)}{\mathbb{E}}_{({{x}},{{y}})\sim\gamma}[\|{{x}}-{{y}}\|^{2}]$
$=\inf_{\gamma\in\Gamma(p,q)}\sum_{i=1}^{n}{\mathbb{E}}_{({{x}},{{y}})\sim\gamma}[({{{x}}^{(i)}}-{{{y}}^{(i)}})^{2}]$
$=\inf_{\gamma\in\Gamma(p,q)}\sum_{i=1}^{n}{\mathbb{E}}_{({{{x}}^{(i)}},{{{y}}^{(i)}})\sim{\gamma^{{(i)}}}}[({{{x}}^{(i)}}-{{{y}}^{(i)}})^{2}],$
where the set of couplings can be rewritten as
$\Gamma(p,q)=\Gamma\left(\prod_{i=1}^{n}{p^{{(i)}}},\prod_{i=1}^{n}{q^{{(i)}}}\right)=\prod_{i=1}^{n}\Gamma({p^{{(i)}}},{q^{{(i)}}}),$
such that
$\inf_{\gamma\in\Gamma(p,q)}\sum_{i=1}^{n}{\mathbb{E}}_{({{{x}}^{(i)}},{{{y}}^{(i)}})\sim{\gamma^{{(i)}}}}[({{{x}}^{(i)}}-{{{y}}^{(i)}})^{2}]$ $=\sum_{i=1}^{n}\inf_{{\gamma^{{(i)}}}\in\Gamma({p^{{(i)}}},{q^{{(i)}}})}{\mathbb{E}}_{({{{x}}^{(i)}},{{{y}}^{(i)}})\sim{\gamma^{{(i)}}}}[({{{x}}^{(i)}}-{{{y}}^{(i)}})^{2}]$
$=\sum_{i=1}^{n}{\mathbb{W}}_{2}^{2}({p^{{(i)}}},{q^{{(i)}}})$
Proposition 24
(Upper Bound on the Wasserstein Distance between Mixtures Distributions, a Generalization of Proposition 4 in Delon and Desolneux (2020)) Let $\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i}\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$ and $\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$ be two mixture distributions, and let $\Gamma({{\pi}}_{p},{{\pi}}_{q})$ be the subset of simplex $\Pi_{M\times N}$ with marginals ${{\pi}}_{p}$ and ${{\pi}}_{q}$. Then for any $\bar{{{\pi}}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q})$, it holds that:
${\mathbb{W}}_{2}^{2}\left(\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i},\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\right)\leq\sum_{i=1}^{M}\sum_{j=1}^{N}{\bar{{{\pi}}}^{{(i,j)}}}{\mathbb{W}}_{2}^{2}(p_{i},q_{j}).$
Proof  Recall that according to the definition of the 2-Wasserstein distance:
${\mathbb{W}}_{2}^{2}\left(\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i},\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\right)=\inf_{\gamma\in\Gamma\left(\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i},\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\right)}{\mathbb{E}}_{({{z}},\tilde{{{z}}})\sim\gamma}[\|{{z}}-\tilde{{{z}}}\|^{2}],$
where $\Gamma\left(\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i},\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\right)$ represents the set of all joint probability distributions with marginal distributions $\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i}$ and $\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}$. Consider the following subset of joint distributions
$\bar{\Gamma}\coloneqq\bigg{\{}\sum_{i=1}^{M}\sum_{j=1}^{N}{\bar{{{\pi}}}^{{(i,j)}}}\gamma\mid\bar{{{\pi}}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q}),\gamma\in\Gamma(p_{i},q_{j})\bigg{\}}\subset\Gamma\left(\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i},\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\right).$
It follows that:
$\inf_{\gamma\in\Gamma\left(\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i},\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}\right)}{\mathbb{E}}_{({{z}},\tilde{{{z}}})\sim\gamma}[\|{{z}}-\tilde{{{z}}}\|^{2}]$ $\leq\inf_{\gamma\in\bar{\Gamma}}{\mathbb{E}}_{({{z}},\tilde{{{z}}})\sim\gamma}[\|{{z}}-\tilde{{{z}}}\|^{2}]$
$=\sum_{i=1}^{M}\sum_{j=1}^{N}{\bar{{{\pi}}}^{{(i,j)}}}\inf_{\gamma\in\Gamma(p_{i},q_{j})}{\mathbb{E}}_{({{z}},\tilde{{{z}}})\sim\gamma}[\|{{z}}-\tilde{{{z}}}\|^{2}]$
$=\sum_{i=1}^{M}\sum_{j=1}^{N}{\bar{{{\pi}}}^{{(i,j)}}}{\mathbb{W}}_{2}^{2}(p_{i},q_{j}),$
which completes the proof.
A.2 Proofs Section 5
Proposition 8: Optimal Transport Plan for Wasserstein Distance Between Distributions and Their Signatures
Proof  The proof follows directly from Theorem 1 in (Ambrogioni et al., 2018). According to this theorem, for any $\rho\in\mathbb{N}$, we have that
${\mathbb{W}}_{\rho}^{\rho}(p,{\Delta}_{\mathcal{C}}\#p)=\sum_{i=1}^{N}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{\rho}\mid{{z}}\in{\mathcal{R}}_{i}].$
Proposition 9: Wasserstein Distance from Signatures of Univariate Gaussians
Proof  From Proposition 8, we have that
${\mathbb{W}}_{2}^{2}({\Delta}_{{\mathcal{C}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))$ $=\sum_{i=1}^{N}\Phi_{i}{\mathbb{E}}_{z\sim\mathcal{N}(0,1)}[({{z}}-c_{i})^{2}\mid z\in[l_{i},u_{i}]].$
where $\Phi_{i}=\mathbb{P}_{z\sim\mathcal{N}(0,1)}[z\in[l_{i},u_{i}]]$. Using the probability density function (pdf) of the Gaussian distribution and a simple change of variables, we can write conditional expectation terms as:

### v2

@@ Appendix A Properties of the Wasserstein Distance
Let $p,q\in{\mathcal{P}}({\mathbb{R}}^{n})$ be joint distributions that are independent across dimensions. Then,
${\mathbb{W}}_{2}^{2}(p,q)=\sum_{l=1}^{n}{\mathbb{W}}_{2}^{2}\left({p^{{(l)}}},{q^{{(l)}}}\right),$ (38)
where $p^{{(l)}}$} and $q^{{(l)}}$} denote the marginals of $p$ and $q$ on the $l$-th dimension.
Proof  See (Panaretos and Zemel, 2019, Section 2, fourth bullet point).
Lemma 31 (${\mathbb{W}}_{2}$ Distance for Mixture Distributions)
Let $p,q\in{\mathcal{P}}({\mathbb{R}}^{n})$ be mixture distributions, i.e., $p=\sum_{i=1}^{M}{{{\pi}}_{p}^{{(i)}}}p_{i}$ and $q=\sum_{j=1}^{N}{{{\pi}}_{q}^{{(j)}}}q_{j}$, where ${{\pi}}_{p}\in\Pi_{M}$, ${{\pi}}_{q}\in\Pi_{N}$, and $p_{i},q_{j}\in{\mathcal{P}}({\mathbb{R}}^{n})$. Then,
${\mathbb{W}}_{2}^{2}(p,q)\leq\inf_{\bar{{{\pi}}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q})}\sum_{i=1}^{M}\sum_{j=1}^{N}{\bar{{{\pi}}}^{{(i,j)}}}{\mathbb{W}}_{2}^{2}(p_{i},q_{j}),$ (39)
where $\Gamma({{\pi}}_{p},{{\pi}}_{q})\subset\Pi_{M\times N}$ denotes the set of couplings between ${{\pi}}_{p}$ and ${{\pi}}_{q}$.
Proof  Consider any coupling $\bar{{\pi}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q})$, and for each pair $(i,j)$ let $\gamma_{i,j}$ be an optimal coupling between $p_{i}$ and $q_{j}$. Then the mixture $\gamma=\sum_{i,j}\bar{\pi}_{i,j}\gamma_{i,j},$ is a valid coupling between $p$ and $q$ with transport cost $\sum_{i,j}\bar{\pi}_{i,j}W_{2}^{2}(p_{i},q_{j})$. Taking the infimum over all $\bar{{{\pi}}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q})$ completes the proof.
Lemma 32 (${\mathbb{W}}_{\rho}$ Distance under Lipschitz Parametrized Stochastic Operations)
Let $\rho\in\mathbb{N}$, $L^{{{w}}}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ be a ${\mathcal{L}}_{L^{{w}}}$-Lipschitz continuous function parametrized by ${{w}}\in{\mathbb{R}}^{d}$, and let w{{w}} be distributed according to $p_{{w}}\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{d})$. Then, for any $p,q\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$, it holds that
${\mathbb{W}}_{\rho}^{\rho}\left({\mathbb{E}}_{{{x}}\sim p}\left[L^{{w}}({{x}})\#p_{{w}}\right],{\mathbb{E}}_{{{z}}\sim q}\left[L^{{w}}({{z}})\#p_{{w}}\right]\right)\leq{\mathbb{E}}_{{{w}}\sim p_{{w}}}\left[{\mathcal{L}}_{L^{{{w}}}}^{\rho}\right]{\mathbb{W}}_{\rho}^{\rho}(p,q).$ (40)
Proof  Since $\|L^{{w}}({{x}})-L^{{w}}({{z}})\|\leq{\mathcal{L}}_{L^{{w}}}\|{{x}}-{{z}}\|$, it follows that
${\mathbb{W}}_{\rho}^{\rho}\left({\mathbb{E}}_{{{x}}\sim p}[L^{{w}}({{x}})\#p_{{w}}],{\mathbb{E}}_{{{z}}\sim q}[L^{{w}}({{z}})\#p_{{w}}]\right)$ $=\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{({{w}},{{x}},{{z}})\sim\gamma}\left[\|L^{{w}}({{x}})-L^{{w}}({{z}})\|^{\rho}\right]$
$\leq\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{({{w}},{{x}},{{z}})\sim\gamma}\left[{\mathcal{L}}_{L^{{w}}}\|{{x}}-{{z}}\|^{\rho}\right].$
Using that $p_{{w}}\times\Gamma(p,q)\subset\Gamma(p_{{w}},p,q)$, we obtain
$\inf_{\gamma\in\Gamma(p_{{w}},p,q)}{\mathbb{E}}_{({{w}},{{x}},{{z}})\sim\gamma}\left[{\mathcal{L}}_{L^{{{w}}}}^{\rho}\|{{x}}-{{z}}\|^{\rho}\right]$ $\leq{\mathbb{E}}_{{{w}}\sim p_{{w}}}\left[{\mathcal{L}}_{L^{{w}}}^{\rho}\right]\inf_{\gamma\in\Gamma(p,q)}{\mathbb{E}}_{({{x}},{{z}})\sim\gamma}\left[\|{{x}}-{{z}}\|^{\rho}\right]$
$={\mathbb{E}}_{{{w}}\sim p_{{w}}}\left[{\mathcal{L}}_{L^{{w}}}^{\rho}\right]{\mathbb{W}}_{\rho}^{\rho}(p,q),$
which concludes the proof.
Appendix B Proofs
We present the proofs for all results discussed in the main text of this work ordered per section.
B.1 Proofs of the Results in Section 5
Proof of Proposition 10: ${\mathbb{W}}_{2}$-Error of the Signature of a Standard Gaussian
From Proposition 9, we have that
${\mathbb{W}}_{2}^{2}({\Delta}_{{\mathcal{C}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))$ $=\sum_{i=1}^{N}{{{\pi}}^{(i)}}{\mathbb{E}}_{z\sim\mathcal{N}(0,1)}\left[({{z}}-c_{i})^{2}\mid z\in[l_{i},u_{i}]\right].$
where
${{{\pi}}^{(i)}}=\mathbb{P}_{z\sim\mathcal{N}(0,1)}[z\in[l_{i},u_{i}]]=\Phi(u_{i})-\Phi(l_{i}).$
Using the closed-form expression for the mean $\mu_{i}{\mathbb{E}}_{z\sim\mathcal{N}(0,1)}[z\mid z\in[l_{i},u_{i}]]$ and variance $\nu_{i}={\mathbb{E}}_{z\sim\mathcal{N}(0,1)}\left[z^{2}\mid z\in[l_{i},u_{i}]\right]-\mu_{i}^{2}$ of a standard univariate Gaussian random variable truncated on $l_{i},u_{i}]$ as given by Eqn (11), respectively, and derived in Section 3 of Manjunath and Wilhelm (2021), we can write the conditional expectation term as:
${\mathbb{E}}_{z\sim\mathcal{N}(0,1)}\left[({{z}}-c_{i})^{2}\mid z\in[l_{i},u_{i}]\right]$
$\qquad={\mathbb{E}}_{z\sim\mathcal{N}(0,1)}\left[{{z}}^{2}\mid z\in[l_{i},u_{i}]\right]-2c_{i}{\mathbb{E}}_{z\sim\mathcal{N}(0,1)}\left[{{z}}\mid z\in[l_{i},u_{i}]\right]+c_{i}^{2}$
$\qquad=\nu_{i}+\mu_{i}^{2}-2c_{i}\mu_{i}+c_{i}^{2}$
$\qquad=\nu_{i}+(\mu_{i}-c_{i})^{2}.$ (41)
Substituting this expression into the original sum concludes the proof.
Proof of Corollary 11: ${\mathbb{W}}_{2}$-Error of the Signature of a Multivariate Gaussian
Let $\tau:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ denote the transformation defined by $\tau({{x}})={{V}}^{T}({{x}}-{{m}})$, where V{{V}} and m{{m}} are, respectively, the (orthogonal) eigenvector matrix and mean of the Gaussian distribution. By Lemma 29, the $2$-Wasserstein is invariant under translations and orthogonal transformations, and thus
${\mathbb{W}}^{2}_{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma))={\mathbb{W}}^{2}_{2}(\tau\#{\Delta}_{\mathcal{C}}\#\mathcal{N}({m},\Sigma),\tau\#\mathcal{N}({m},\Sigma)).$
Since, $\Sigma={{V}}\mathrm{diag}({{\lambda}}){{V}}^{T}$, it holds that
$\tau\#\mathcal{N}({m},\Sigma)=\mathcal{N}({{\bar{0}}},\mathrm{diag}({{\lambda}}))=\mathrm{diag}({{\lambda}})^{1/2}\#\mathcal{N}({{\bar{0}}},{{I}}),$
where, with a slight abuse of notation, we write $\mathrm{diag}({{\lambda}})^{1/2}\#p$ to denote the pushforward under the linear map ${{x}}\mapsto\mathrm{diag}({{\lambda}})^{1/2}{{x}}$. Next, consider the pushforward of the signature:
$\tau\#{\Delta}_{\mathcal{C}}\#\mathcal{N}({m},\Sigma)$ $={\Delta}_{\mathrm{Post}({\mathcal{C}}-\{{{m}}\},{{V}}^{T})}\#\tau\#\mathcal{N}({m},\Sigma)$
$={\Delta}_{\mathrm{Post}({\mathcal{C}}-\{{{m}}\},{{V}}^{T})}\#\mathrm{diag}(\lambda)^{1/2}\#\mathcal{N}({{\bar{0}}},{{I}})$
$=\mathrm{diag}(\lambda)^{1/2}\#{\Delta}_{\mathrm{Post}({\mathcal{C}}-\{{{m}}\},{{T}})}\#\mathcal{N}({{\bar{0}}},{{I}})$
$=\mathrm{diag}(\lambda)^{1/2}\#{\Delta}_{{\mathcal{C}}^{1}\times{\mathcal{C}}^{2}\times\ldots\times{\mathcal{C}}^{n}}\#\mathcal{N}({{\bar{0}}},{{I}}),$
where the final equality holds due to the axis-aligned grid assumption in (12). Since both terms are product distributions, we can apply Lemma 30, yielding
${\mathbb{W}}^{2}_{2}(\tau\#{\Delta}_{\mathcal{C}}\#\mathcal{N}({m},\Sigma),\tau\#\mathcal{N}({m},\Sigma))$ $=\sum_{l=1}^{n}{\mathbb{W}}_{2}^{2}\left(\sqrt{{{{\lambda}}^{(l)}}}\#\mathcal{N}(0,1),\sqrt{{{{\lambda}}^{(l)}}}\#{\Delta}_{{\mathcal{C}}^{l}}\#\mathcal{N}(0,1)\right)$
$=\sum_{l=1}^{n}{\mathbb{W}}_{2}^{2}{{{\lambda}}^{(l)}}\left(\mathcal{N}(0,1),{\Delta}_{{\mathcal{C}}^{l}}\#\mathcal{N}(0,1)\right),$
where the final equality follows from the homogeneity of the $2$-Wasserstein distance. This completes the proof.
Proof of Proposition 13:
If $N=1$, Eqn (15) gives ${\mathcal{C}}=\{0\}$ and Proposition 9 yields
${\mathbb{W}}_{2}^{2}({\Delta}_{\{0\}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))=\int z^{2}\,\phi(dz)=1=\frac{4\log 1}{1^{2}}+r(1)=r(1),$
since $\log 1=0$ and $r(1)=1$ by Eqn (17), and where $\phi$ is the pdf of a standard (univariate) Gaussian distribution. Hence the bound holds for $N=1$. For the remainder of the proof assume $N\geq 2$. Let $\eta=\frac{2\sqrt{\log N}}{N}$ and ${\mathcal{C}}=\{c_{i}\}_{i=1}^{N}$ with $c_{i}=\eta(2i-N-1)$. The induced partition (Eqn (7)) is
${\mathcal{R}}_{1}=\left[-\infty,c_{1}+\eta\right],\quad{\mathcal{R}}_{i}=\left[c_{i}-\eta,c_{i}+\eta\right],\quad\forall i\in\{2,\ldots,N-1\},\quad{\mathcal{R}}_{N}=\left[c_{N}-\eta,\infty\right].$
By Proposition 9, we have
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))$
$\qquad=\sum_{i=2}^{N-1}\int_{c_{i}-\eta}^{c_{i}+\eta}(z-c_{i})^{2}\phi(dz)+\int_{-\infty}^{c_{1}+\eta}(z-c_{1})^{2}\phi(dz)+\int_{c_{N}-\eta}^{\infty}(z-c_{N})^{2}\phi(dz)$
$\qquad=\sum_{i=1}^{N}\int_{c_{i}-\eta}^{c_{i}+\eta}(z-c_{i})^{2}\phi(dz)+\int_{-\infty}^{-\eta N}(z-c_{1})^{2}\phi(dz)+\int_{\eta N}^{\infty}(z-c_{N})^{2}\phi(dz),$
where in the last step we split the unbounded regions and used $c_{1}=-\eta N+\eta$, $c_{N}=\eta N-\eta$.
Bounded integrals: since $|z-c_{i}|\leq\eta$ for all $z\in{\mathcal{R}}_{i}$,
$\sum_{i=1}^{N}\int_{c_{i}-\eta}^{c_{i}+\eta}(z-c_{i})^{2}\phi(dz)$ $\leq\eta^{2}\sum_{i=1}^{N}[\Phi(c_{i}+\eta)-\Phi(c_{i}-\eta)]$
$=\eta^{2}\left[\Phi\left(\eta N\right)-\Phi\left(-\eta N\right)\right]=\eta^{2}\left[2\Phi\left(\eta N\right)-1\right],$
where the last equalities follow from symmetry and telescoping of the CDF terms.
Unbounded integrals: by symmetry ($c_{1}=-c_{N}$),
$\int_{-\infty}^{-\eta N}(z-c_{1})^{2}\phi(dz)+\int_{\eta N}^{\infty}(z-c_{N})^{2}\phi(dz)=2\int_{\eta N}^{\infty}(z-c_{N})^{2}\phi(dz).$
For $z\geq\eta N$, $z-c_{N}=z-\eta N+\eta\leq z$, hence
$\int_{\eta N}^{\infty}(z-c_{N})^{2}\mathcal{N}(dz\mid 0,1)\leq\int_{\eta N}^{\infty}z^{2}\mathcal{N}(dz\mid 0,1)=[1-\Phi(\eta N)]+\eta N\phi(\eta N),$
using the truncated second-moment identity for standard Gaussians (Eqn (41)).
Combining and denoting $L=\eta N=2\sqrt{\log N}$,
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))$ $\leq\eta^{2}\left[2\Phi\left(L\right)-1\right]+2[1-\Phi(L)]+2L\phi(L)$
$=-2\eta^{2}\left[1-\Phi\left(L\right)\right]+\eta^{2}+2[1-\Phi(L)]+2L\phi(L)$
$=2(1-\eta^{2})\left(1-\Phi\left(L\right)\right)+\eta^{2}+2L\phi\left(L\right)$
$=2\left(1-\frac{L^{2}}{N^{2}}\right)\left(1-\Phi\left(L\right)\right)+\frac{L^{2}}{N^{2}}+2L\phi\left(L\right)$ (*)
By Mills’ ratio $1-\Phi(L)\leq\frac{\phi(L)}{L}$ and $0\leq\frac{L^{2}}{N^{2}}\leq 1$, we bound the first term in ()(*) by $2\frac{\phi(L)}{L}$, yielding
$(*)$ $\leq 2\frac{\phi(L)}{L}+\frac{L^{2}}{N^{2}}+2L\phi(L)=2\phi(L)\left(L+\frac{1}{L}\right)+\frac{L^{2}}{N^{2}}.$
With $L=2\sqrt{\log N}$ we have $\phi(L)=\frac{1}{\sqrt{2\pi}}e^{-L^{2}/2}=\frac{1}{\sqrt{2\pi}}e^{-2\log N}=\frac{1}{N^{2}\sqrt{2\pi}}$, so
$2\phi(L)\left(L+\tfrac{1}{L}\right)=\frac{4}{N^{2}\sqrt{2\pi}}\left(\sqrt{\log N}+\frac{1}{4\sqrt{\log N}}\right).$
Therefore
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))\leq\frac{4\log N}{N^{2}}+\frac{4}{N^{2}\sqrt{2\pi}}\left(\sqrt{\log N}+\frac{1}{4\sqrt{\log N}}\right)=\frac{4\log N}{N^{2}}+r(N).$
This completes the proof.
Proof of Corollary 15
From Eqn (14), the $2$-Wasserstein error of the component-wise signature of a Gaussian mixture is bounded by the mixture-weighted sum of the signature errors of its components:
${\mathbb{W}}_{2}^{2}\big{(}p,\bm{{\Delta}}_{\{{\mathcal{C}}_{i}\}_{i=1}^{M}}\#p\big{)}\leq\sum_{i=1}^{M}\tilde{\pi}_{i}\,{\mathbb{W}}_{2}^{2}\big{(}\mathcal{N}(m_{i},\Sigma_{i}),{\Delta}_{{\mathcal{C}}_{i}}\#\mathcal{N}(m_{i},\Sigma_{i})\big{)}.$
For each component $i$, diagonalize $\Sigma_{i}={{V}}_{i}\mathrm{diag}({{\lambda}}_{i}){{V}}_{i}^{T}$ and apply Corollary 11 together with the axis-aligned grid assumption (12) to obtain
${\mathbb{W}}_{2}^{2}\big{(}\mathcal{N}(m_{i},\Sigma_{i}),{\Delta}_{{\mathcal{C}}_{i}}\#\mathcal{N}(m_{i},\Sigma_{i})\big{)}=\sum_{j=1}^{n}\lambda_{i,j}\,{\mathbb{W}}_{2}^{2}\big{(}\mathcal{N}(0,1),{\Delta}_{{\mathcal{C}}_{i}^{j}}\#\mathcal{N}(0,1)\big{)}.$
Proposition 13 applied to each one-dimensional uniform grid of points ${\mathcal{C}}_{i}^{j}$ (of size $N_{i,j}$) yields
${\mathbb{W}}_{2}^{2}\big{(}\mathcal{N}(0,1),{\Delta}_{{\mathcal{C}}_{i}^{j}}\#\mathcal{N}(0,1)\big{)}\leq\frac{4\log N_{i,j}}{N_{i,j}^{2}}+r(N_{i,j}),$
with $r(\cdot)$ given in Eqn (17). Combining the three displays finishes the proof: